1. Find the matrix A of the reflection Refl in Rw.r.t. line : +2y = 0. Is A invertible? If so, find the inverse

could you please provide all the steps, I'm not sure how to find the reflection or how to know if it is invertable.
Answer #1

line Li atzy y = (-1/2) x (tono) a comparing with = ma (or t Wc get tano = -1 2 MILO Cos20 = 1 tan²o It tan o 1 + 114 Sin20 =